Please help me find the standard form of the equation of the line that is perpendicular to 6x+4y=-2 and contains (-2,-4)
A. 2x-3y=8
B.x+2y=-8
C.3x+2y=-8
D.2x+3y=-8

Answer by jim_thompson5910(35256)

Anything perpendicular to 6x+4y=-2 will be of the form 4x-6y=C where C is any number


Since this perpendicular line goes through (-2,-4), we know that x=-2 and y=-4. Use this to find C


4x-6y=C


4(-2)-6(-4)=C


-8+24=C


16=C


C = 16


So the equation is 4x-6y=16 and you can divide every term by 2 to get 2x-3y=8


So the answer is choice A